# Physical synchronization of soft self-oscillating limbs for fast and autonomous locomotion

### v1

@@ Acknowledgments
We thank all members of our Soft Robotic Matter Group for the invaluable discussions. We thank Corentin Coulais and Said RK Rodríguez for the constructive feedback. Funding: A.C., H.A.H.S., and J.T.B.O. acknowledge the European Union’s 2020 ERC-STG under grant agreement No. 948132. This work is part of the Dutch Research Council (NWO) and was performed at the research institute AMOLF. Authors contributions: A.C. and J.T.B.O. proposed and developed the research idea; A.C. designed and fabricated the devices; A.C. and H.A.H.S. performed the experiments; A.C. and H.A.H.S. performed the data analysis; A.C. and J.T.B.O. wrote the manuscript; A.C., H.A.H.S., and J.T.B.O. revised the manuscript; J.T.B.O. supervised the research. Competing interests: A.C., H.A.H.S., and J.T.B.O. declare no competing interests.

### v2

@@ Acknowledgments
We thank all members of our Soft Robotic Matter Group for the invaluable discussions. We thank Niels Commandeur for technical support, Sergio Picella for the input on the analytical model, and Corentin Coulais and Said RK Rodríguez for the constructive feedback. Funding: J.T.B.O. acknowledges the European Union’s 2020 ERC-STG under grant agreement No. 948132. This work is part of the Dutch Research Council (NWO) and was performed at the research institute AMOLF. Authors contributions: A.C. and J.T.B.O. proposed and developed the research idea; A.C. designed and fabricated the devices; A.C. and H.A.H.S. performed the experiments; A.C. and H.A.H.S. performed the data analysis; A.C. created the figures and videos; A.C. and J.T.B.O. wrote the manuscript; A.C., H.A.H.S., and J.T.B.O. revised the manuscript; J.T.B.O. supervised the research. Competing interests: A.C., H.A.H.S., and J.T.B.O. declare no competing interests. Data and materials availability: All data is available on Zenodo (?).
Supplementary Materials
The PDF file includes:
Materials and Methods
Supplementary Text
Figs. S1 to S27
Tables S1 to S4
Captions for Movies S1 to S7
References (43-62)
Other Supplementary Material for this manuscript includes the following:
Movies S1 to S7